OpenStudy (anonymous):

Simplify 6x^4-9x^2+12x/3x help pleasee

OpenStudy (agent0smith):

This may help...\[\Large \frac{ 6x^4-9x^2+12x }{ 3x} = \] \[\Large \frac{ 6x^4 }{ 3x} - \frac{9x^2 }{ 3x} + \frac{12x }{ 3x}\] Now you can try to simplify each piece.

OpenStudy (anonymous):

Wild it be 2x^3-3x+4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Yes it would! :)

OpenStudy (anonymous):

Can you help me with mother one ? Pleasssseee?

OpenStudy (anonymous):

Another ""

OpenStudy (uri):

Wow mother one....

OpenStudy (uri):

Close this ask new question in a new thread.